OpenStudy (anonymous):

solve by completing the square. 4x^2-29x+7=0??

OpenStudy (anonymous):

(4x-1)(x-7)=0 s.t x1=1/4,x2=7

OpenStudy (saifoo.khan):

@strawberry_militia , do you know how to do completing square?

OpenStudy (anonymous):

erm sort of my teacher said to put (ax+b)+c and equate terms.

OpenStudy (anonymous):

e.g a^2+x^2=4x^2???

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=gzm-uhj06q8

OpenStudy (anonymous):

khan academy????

OpenStudy (saifoo.khan):

Exactly.

OpenStudy (anonymous):

he completes the square differently to my teacher.

OpenStudy (saifoo.khan):

Atleast we can understand his method easily. ;)